(D^2 -4DD^'+ 3D^'^2)z =0


1
Expert's answer
2021-01-13T16:50:22-0500

Rewrite this equation in the next view:

"z_{xx}-4z_{xy}+3z_{yy}=0"

Now lead it to canonical view. To do this, let input new variables:

"\begin{cases}
\eta=y+x\\
\xi= y+3x\\
\end{cases}"

So now we have:

"\dfrac{\partial z}{\partial x}=\dfrac{\partial z}{\partial \eta}+3\dfrac{\partial z}{\partial \xi}"

"\dfrac{\partial z}{\partial y}=\dfrac{\partial z}{\partial \eta}+\dfrac{\partial z}{\partial \xi}"

"\dfrac{\partial^2 z}{\partial x^2}=(\dfrac{\partial z}{\partial \eta}+3\dfrac{\partial z}{\partial \xi})^2"

"\dfrac{\partial^2 z}{\partial y^2}=(\dfrac{\partial z}{\partial \eta}+\dfrac{\partial z}{\partial \xi})^2"

"\dfrac{\partial^2 z}{\partial x \partial y}=\dfrac{\partial z^2}{\partial \eta^2}+4\dfrac{\partial^2 z}{\partial \xi \partial\eta} + 9\dfrac{\partial z^2}{\partial \xi^2}"

"z_{\eta\eta}+6z_{\eta\xi}+9z_{\xi\xi}-4(z_{\eta\eta}+4z_{\eta\xi} +9z_{\xi\xi})+3(z_{\eta\eta}+2z_{\eta\xi}+z_{\xi\xi})=0"

The canonical equation:

"6z_{\xi\xi}+z_{\xi\eta}=0"

"\dfrac{\partial}{\partial \xi}(6\dfrac{\partial z}{\partial \xi} + \dfrac{\partial z}{\partial \eta})=0"

"6\dfrac{\partial z}{\partial \xi} + \dfrac{\partial z}{\partial \eta}=C_1(\eta)"

"\dfrac{\partial \xi}{6} =\partial \eta=\dfrac{\partial z}{C_1(\eta)}"

"z_1=\int C_1(\eta) d\eta +C_2"

"z_2=C_1(y+x)*(y+3x)+C_3"